### v1

# Completeness of coalition logics with seriality, independence of agents, or determinism

Abstract
Coalition Logic is a central logic in logical research on strategic reasoning. In a recent paper, Li and Ju argued that generally, models of Coalition Logic, concurrent game models, have three too strong assumptions: seriality, independence of agents, and determinism. They presented a Minimal Coalition Logic based on general concurrent game models, which do not have the three assumptions. However, when constructing coalition logics about strategic reasoning in special kinds of situations, we may want to keep some of the assumptions. Thus, studying coalition logics with some of these assumptions makes good sense. In this paper, we show the completeness of these coalition logics by a uniform approach.
Keywords: coalition logics, general concurrent game models, seriality, independence of agents, determinism, completeness
1 Introduction
1.1 Coalition Logic
Coalition Logic $\mathsf{CL}$ ([7, 8]) is a central logic in logical research on strategic reasoning. Many important logics about strategic reasoning are extensions of $\mathsf{CL}$: Alternating-time Temporal Logic $\mathsf{ATL}$ ([1]) is a temporal extension of $\mathsf{CL}$; Strategy Logic $\mathsf{SL}$ [6] is an extension of $\mathsf{CL}$ whose language has quantifiers for and names of strategies; and so on. We refer to [2], [9], and [11] for overviews of the area.
The language of $\mathsf{CL}$ is a modal language with the featured operator $\langle[\mathrm{A}]\rangle\phi$, indicating some available joint action of the coalition $\mathrm{A}$ ensures $\phi$. Its models are concurrent game models. Roughly, in a concurrent game model: there are some states; there are some agents, who can form coalitions; at every state, every coalition has some available joint actions; every joint action of a coalition has some possible outcome states. The formula $\langle[\mathrm{A}]\rangle\phi$ is true at a state in a concurrent game model if $\mathrm{A}$ has an available joint action such that $\phi$ is true at every possible outcome state of the action.
1.2 Minimal Coalition Logic
Li and Ju [5] argued that generally, concurrent game models have three too strong assumptions. The first one is seriality: coalitions always have available joint actions. The second one is the independence of agents: the merge of two available joint actions of two disjoint coalitions is always an available joint action of the union of the two coalitions. The third one is determinism: joint actions of the grand coalition always have a unique outcome. Based on general concurrent game models, which do not have the three assumptions, Li and Ju [5] presented a Minimal Coalition Logic $\mathsf{MCL}$.
1.3 It makes good sense to study the logics between Minimal Coalition Logic and Coalition Logic
Although the three assumptions are generally too strong, we might want to keep some of them when constructing logics for strategic reasoning in special kinds of situations. First, consider seriality. When assuming that agents can always perform the special action doing nothing, seriality holds. Second, consider the independence of agents. In many situations, agents’ actions do not interfere with each other, and the independence of agents holds. Consider the game of rock-paper-scissors and the Prisoner’s Dilemma as examples. As mentioned by Sergot [10], when we just consider some special kind of actions, such as attempting to do something, the independence of agents does hold. Third, consider determinism. Typical games are designed such that no matter how players behave, there is always a deterministic outcome. Again, consider the game of rock-paper-scissors as an example. In addition, when assuming that nature is a player, determinism holds, at least on the macro level.
1.4 Our work
According to which of the three properties we want to keep, there are eight coalition logics. The completeness $\mathsf{CL}$ is proved by Pauly [8], and the completeness of $\mathsf{MCL}$ is shown by Li and Ju [5]. The completeness of the other six logics has not been shown yet.
In this work, we show the completeness of the eight logics by an uniform approach. Generally speaking, our proof approach is as follows. First, we transform every formula to a conjunction of some standard disjunctions. Second, we show a downward validity lemma, which reduces the validity of a standard disjunction to the validity of some formulas with lower modal depth. Third, we show an upward derivability lemma, which reduces the derivability of a standard disjunction to the derivability of some formulas with lower modal depth. Fourth, we show the completeness by induction.
Generally, our approach follows the approach of Goranko and van Drimmelen [3], who showed the completeness of $\mathsf{ATL}$, and Li and Ju [5], although there is much difference between them in detail. Pauly [8] showed the completeness of $\mathsf{CL}$ by the following approach: first, he showed a representation theorem and transformed the semantics for $\mathsf{CL}$ based on concurrent game models to a kind of neighborhood semantics; second, he showed the completeness of $\mathsf{CL}$ with respect to the neighborhood semantics by the Henkin method222 Goranko, Jamroga, and Turrini [4] pointed out that the representation theorem given by Pauly [8] is incorrect, and consequently, the neighborhood semantics given by Pauly [8] is incorrect. They offered the correct representation theorem and neighborhood semantics. However, they showed that the neighborhood semantics given by Pauly [8] determines the same set of valid formulas as the correct neighborhood semantics. . Our approach is different from this approach.
The rest of the paper is organized as follows. In Section 2, we present the eight coalition logics, including their models, language, semantics, and axiomatic systems. In Section 3, we show the completeness of the eight logics. We point out some further work in Section 4.
2 Eight coalition logics determined by eight classes of general concurrent game models
2.1 General concurrent game models and three properties
Let $\mathtt{AG}$ be a finite set of agents. Subsets of $\mathtt{AG}$ are called coalitions. $\mathtt{AG}$ is called the grand coalition. Let $\mathtt{AC}$ be a nonempty set of actions. For every coalition $\mathrm{C}$, a function ${\sigma_{\mathrm{C}}}:\mathrm{C}\to\mathtt{AC}$ is called a joint action of $\mathrm{C}$. The joint actions of $\mathtt{AG}$ are called action profiles. For every $\mathrm{C}\subseteq\mathtt{AG}$, we define $\mathtt{JA}_{\mathrm{C}}=\{\sigma_{\mathrm{C}}\mid\sigma_{\mathrm{C}}:\mathrm{C}\rightarrow\mathtt{AC}\}$, which is the set of joint actions of $\mathrm{C}$. We define $\mathtt{JA}=\bigcup\{\mathtt{JA}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$.
Definition 1 (General concurrent game frames).
A general concurrent game frame is a tuple $\mathfrak{F}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}})$, where:
  • $\mathtt{ST}$ is a nonempty set of states;
  • $\mathtt{AG}$ is a finite set of agents;
  • $\mathtt{AC}$ is a nonempty set of actions;
  • $\mathsf{out}_{\mathtt{AG}}:\mathtt{ST}\times\mathtt{JA}_{\mathtt{AG}}\rightarrow\mathcal{P}(\mathtt{ST})$ is an outcome function for $\mathtt{AG}$.
General concurrent game frames will be simply called frames in the sequel.
Definition 2 (Outcome functions and availability functions).
Let $\mathfrak{F}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}})$ be a frame.
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the outcome function $\mathsf{out}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$ and ${\sigma_{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$,
$\mathsf{out}_{\mathrm{C}}(s,{\sigma_{\mathrm{C}}})=\bigcup\{\mathsf{out}_{\mathtt{AG}}(s,{\sigma_{\mathtt{AG}}})\mid{\sigma_{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }{\sigma_{\mathrm{C}}}\subseteq{\sigma_{\mathtt{AG}}}\}.$
For every $\mathrm{C}\subseteq\mathtt{AG}$, define the availability function $\mathtt{aja}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$:
$\mathtt{aja}_{\mathrm{C}}(s)=\{{\sigma_{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathsf{out}_{\mathrm{C}}(s,{\sigma_{\mathrm{C}}})\neq\emptyset\}.$
The elements of ${\mathtt{aja}(s,\mathrm{C})}$ are called available joint actions of $\mathrm{C}$ at $s$.
Definition 3 (Three properties of frames).
Let $\mathfrak{F}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}})$ be a frame. We say:
  • $\mathfrak{F}$ is serial if $\mathtt{aja}_{\mathrm{C}}(s)\neq\emptyset$ for all $s\in\mathtt{ST}$ and $\mathrm{C}\subseteq\mathtt{AG}$;
  • $\mathfrak{F}$ is independent if for all $s\in\mathtt{ST}$, $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$, ${\sigma_{\mathrm{C}}}\in\mathtt{aja}_{\mathrm{C}}(s)$, and ${\sigma_{\mathrm{D}}}\in\mathtt{aja}_{\mathrm{D}}(s)$, if $\mathrm{C}\cap\mathrm{D}=\emptyset$, then ${\sigma_{\mathrm{C}}}\cup{\sigma_{\mathrm{D}}}\in\mathtt{aja}_{\mathrm{C}\cup\mathrm{D}}(s)$;
  • $\mathfrak{F}$ is deterministic if for all $s\in\mathtt{ST}$ and ${\sigma_{\mathtt{AG}}}\in\mathtt{aja}_{\mathtt{AG}}(s)$, ${\mathsf{out}_{\mathtt{AG}}(s,{\sigma_{\mathtt{AG}}})}$ is a singleton.
We let the three properties correspond to the symbols $\mathsf{S}$, $\mathsf{I}$, and $\mathsf{D}$, respectively. We let the eight combinations of the three properties correspond to the eight strings $\epsilon$, $\mathsf{S}$, $\mathsf{I}$, $\mathsf{D}$, $\mathsf{S}\mathsf{I}$, $\mathsf{S}\mathsf{D}$, $\mathsf{I}\mathsf{D}$, and $\mathsf{S}\mathsf{I}\mathsf{D}$, respectively. We use $\mathrm{ES}$ to indicate the set of the eight strings.
For any $\mathsf{X}\in\mathrm{ES}$, we say a frame $\mathfrak{F}$ is an $\mathsf{X}$-frame if $\mathfrak{F}$ has the properties corresponding to $\mathsf{X}$.
Definition 4 (General concurrent game models).
A general concurrent game model is a tuple $\mathtt{M}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathtt{AP},\mathsf{label})$, where:
  • $(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}})$ is a frame;
  • $\mathsf{AP}$ is a countable set of atomic propositions;
  • $\mathsf{label}:\mathtt{ST}\rightarrow\mathcal{P}(\mathtt{AP})$ is a labeling function.
General concurrent game models will simply be called models in the sequel. For any model $\mathtt{M}$ and state $s$ of $\mathtt{M}$, $(\mathtt{M},s)$ is called a pointed model.
For any $\mathsf{X}\in\mathrm{ES}$, a model $\mathtt{M}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathtt{AP},\mathsf{label})$ is called an $\mathsf{X}$-model if $(\mathtt{ST},\mathtt{AG},\mathtt{AC},$ $\mathsf{out}_{\mathtt{AG}})$ is an $\mathsf{X}$-frame. Note that $\epsilon$-models are models of $\mathsf{MCL}$ and $\mathsf{S}\mathsf{I}\mathsf{D}$-models are models of $\mathsf{CL}$.
2.2 Language
Definition 5 (The language $\Phi$).
The language $\Phi$ is defined as follows, where $p$ ranges over $\mathtt{AP}$ and $\mathrm{C}\subseteq\mathtt{AG}$:
$\phi::=\top\mid p\mid\neg\phi\mid(\phi\wedge\phi)\mid\langle[\mathrm{C}]\rangle\phi$
The formula $\langle[\mathrm{C}]\rangle\phi$ indicates that some available joint action of $\mathrm{C}$ ensures $\phi$. Define the propositional connectives ,$\bot$,$\lor$,$\rightarrow$ , ∨ , → and $\leftrightarrow$ as usual. Define the dual $[\langle\mathrm{C}\rangle]\phi$ of $\langle[\mathrm{C}]\rangle\phi$ as $\neg\langle[\mathrm{C}]\rangle\neg\phi$, indicating every available joint action of $\mathrm{A}$ enables $\phi$.
2.3 Semantics
Definition 6 (Semantics of $\Phi$).
Let $\mathtt{M}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathtt{AP},\mathsf{label})$ be a model.
$\mathtt{M},s\Vdash\top$ always
$\mathtt{M},s\Vdash p$Leftrightarrow $p\in\mathsf{label}(s)$
$\mathtt{M},s\Vdash\neg\phi$Leftrightarrow not $\mathtt{M},s\Vdash\phi$
$\mathtt{M},s\Vdash\phi\land\psi$Leftrightarrow $\mathtt{M},s\Vdash\phi$ and $\mathtt{M},s\Vdash\psi$
$\mathtt{M},s\Vdash\langle[\mathrm{C}]\rangle\phi$Leftrightarrow there is ${\sigma_{\mathrm{C}}}\in\mathtt{aja}_{\mathrm{C}}(s)$ such that for all $t\in\mathsf{out}_{\mathrm{C}}(s,{\sigma_{\mathrm{C}}})$, $\mathtt{M},t\Vdash\phi$
It can be verified:
$\mathtt{M},s\Vdash[\langle\mathrm{C}\rangle]\phi$Leftrightarrow for all $\sigma_{\mathrm{C}}\in\mathtt{aja}_{\mathrm{C}}(s)$, there is $t\in\mathsf{out}_{\mathrm{C}}(s,\sigma_{\mathrm{C}})$ such that $\mathtt{M},t\Vdash\phi$
For any $\mathsf{X}\in\mathrm{ES}$, set of formulas $\Gamma$, formula $\phi$, we define the following notions in the expected way: $\phi$ is $\mathsf{X}$-valid ($\models_{\mathsf{MCL}+\mathsf{X}}\phi$), $\phi$ is $\mathsf{X}$-satisfiable, and $\phi$ is a logical consequence of $\Gamma$ ($\Gamma\models_{\mathsf{MCL}+\mathsf{X}}\phi)$.
Let $(\mathtt{M},s)$ be a pointed model and ${\sigma_{\mathrm{C}}}$ be a (possibly unavailable) joint action at $s$. We use ${\sigma_{\mathrm{C}}}\vartriangleright_{(\mathtt{M},s)}\phi$ to denote ${\sigma_{\mathrm{C}}}$ ensures $\phi$ at $(\mathtt{M},s)$, that is, $\mathtt{M},t\Vdash\phi$ for all $t\in\mathsf{out}_{\mathrm{C}}(s,{\sigma_{\mathrm{C}}})$. We use ${\sigma_{\mathrm{C}}}\ltimes_{(\mathtt{M},s)}\phi$ to denote that ${\sigma_{\mathrm{C}}}$ enables $\phi$ at $(\mathtt{M},s)$, that is, $\mathtt{M},t\Vdash\phi$ for some $t\in\mathsf{out}_{\mathrm{C}}(s,{\sigma_{\mathrm{C}}})$.
2.4 Axiomatic systems
We let the following formulas respectively correspond to the symbols $\mathsf{S}$, $\mathsf{I}$ and $\mathsf{D}$:
Seriality ($\mathtt{A}\text{-}\mathtt{Ser}$): $\langle[\mathrm{C}]\rangle\top$
Independence of agents ($\mathtt{A}\text{-}\mathtt{IA}$): $(\langle[\mathrm{C}]\rangle\phi\land\langle[\mathrm{D}]\rangle\psi)\rightarrow\langle[\mathrm{C}\cup\mathrm{D}]\rangle(\phi\land\psi)$, where $\mathrm{C}\cap\mathrm{D}=\emptyset$
Determinism ($\mathtt{A}\text{-}\mathtt{Det}$): $\langle[\mathrm{C}]\rangle(\phi\lor\psi)\rightarrow(\langle[\mathrm{C}]\rangle\phi\lor\langle[\mathtt{AG}]\rangle\psi)$
Definition 7 (Axiomatic systems for ${\mathsf{MCL}+\mathsf{X}}$).
For all $\mathsf{X}$ in $\mathrm{ES}$, the axiomatic system for ${\mathsf{MCL}+\mathsf{X}}$ consists of the following axioms and inference rules:
Axioms:
Tautologies ($\mathtt{A}\text{-}\mathtt{Tau}$): all propositional tautologies
No absurd available actions ($\mathtt{A}\text{-}\mathtt{NAAA}$): $\neg\langle[\mathrm{C}]\rangle\bot$
Monotonicity of goals ($\mathtt{A}\text{-}\mathtt{MG}$): $\langle[\emptyset]\rangle(\phi\rightarrow\psi)\rightarrow(\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\mathrm{C}]\rangle\psi)$
Monotonicity of coalitions ($\mathtt{A}\text{-}\mathtt{MC}$): $\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\mathrm{D}]\rangle\phi$, where $\mathrm{C}\subseteq\mathrm{D}$
$\mathtt{A}\text{-}\mathsf{X}$: the formulas corresponding to the elements of $\mathsf{X}$
Inference rules:
Modus ponens ($\mathtt{MP}$): $\dfrac{\phi,\phi\rightarrow\psi}{\psi}$
Conditional necessitation ($\mathtt{CN}$): $\dfrac{\phi}{\langle[\mathrm{C}]\rangle\psi\rightarrow\langle[\emptyset]\rangle\phi}$
For any $\mathsf{X}$ in $\mathrm{ES}$, we use $\vdash_{\mathsf{MCL}+\mathsf{X}}\phi$ to indicate $\phi$ is derivable in the system for ${\mathsf{MCL}+\mathsf{X}}$.
The following fact will be used later, whose proof is skipped:
Fact 1.
The following formulas and rule are derivable in $\mathsf{MCL}$:
Condition for empty action ($\mathtt{A}\text{-}\mathtt{CEA}$): $\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\emptyset]\rangle\top$
Special independence of agents ($\mathtt{A}\text{-}\mathtt{SIA}$): $(\langle[\emptyset]\rangle\phi\land\langle[\mathrm{C}]\rangle\psi)\rightarrow\langle[\mathrm{C}]\rangle(\phi\land\psi)$
Monotonicity ($\mathtt{Mon}$): from $\phi\rightarrow\psi$, we can get $\langle[\mathrm{C}]\rangle\phi\rightarrow\langle[\mathrm{D}]\rangle\psi$,
where $\mathrm{C}\subseteq\mathrm{D}$.
Remarks
The axiomatic system for $\mathsf{MCL}$ given in Definition 7 is shown to be sound and complete with respect to the set of $\epsilon$-valid formulas in [5]. By [5], the axiomatic system for $\mathsf{CL}$ given in Definition 7 is equivalent to the sound and complete axiomatic system for $\mathsf{CL}$ given in [8].
The formula $\langle[\mathrm{C}]\rangle\top$ defines the class of frames where $\mathrm{C}$ at every state has an available joint action. However, the formula $(\langle[\mathrm{C}]\rangle p\land\langle[\mathrm{D}]\rangle q)\rightarrow\langle[\mathrm{C}\cup\mathrm{D}]\rangle(p\land q)$, where $\mathrm{C}$ and $\mathrm{D}$ are disjoint, does not define the class of frames where at every state, the union of any joint actions of $\mathrm{C}$ and $\mathrm{D}$ is a joint action of $\mathrm{C}\cup\mathrm{D}$; the formula $\langle[\mathrm{C}]\rangle(p\lor q)\rightarrow(\langle[\mathrm{C}]\rangle p\lor\langle[\mathtt{AG}]\rangle q)$ does not define the class of frames where at every state, every available joint action of $\mathtt{AG}$ has a unique outcome state.
In the literature, the formula ¬[]$\neg\langle[\emptyset]\rangle\neg\phi\rightarrow\langle[\mathtt{AG}]\rangle\phi$ is commonly treated as an axiom related to determinism. We want to point out that it is not valid in all deterministic frames, although it is valid in all serial and deterministic frames.
3 Completeness of the eight coalition logics
3.1 Our approach
To show the completeness of an axiomatic system, it suffices to show that every valid formula is derivable in it. We will achieve this by induction on the modal depth of formulas.
The basic step of the induction can be directly done by the classical propositional logic, and the inductive step is crucial. We will go through the inductive step in the following way.
First, we transfer every formula to a normal form, which is a conjunction of some standard disjunctions, that are disjunctions meeting some conditions.
Second, we show the downward validity lemma: for every standard disjunction $\phi$, if $\phi$ is valid, then the validity-reduction-condition of $\phi$ is met. Here the validity-reduction-condition of $\phi$ concerns the validity of some formulas with lower modal depth than $\phi$.
Third, we show the upward derivability lemma: for every standard disjunction $\phi$, if the derivability-reduction-condition of $\phi$ is met, then $\phi$ is derivable. Here, the derivability-reduction-condition of $\phi$ is the result of replacing validity in the validity-reduction-condition of $\phi$ by derivability.
Note a conjunction is valid (derivable) if and only if all of its conjuncts are valid (derivable). Thus, the two lemmas would let us go through the inductive step. Note by soundness and completeness, the other direction of the two lemmas also holds, which is why their names contain “reduction”.
The downward validity lemma is the most crucial step for the whole proof. We will show it by showing its contrapositive, called the upward satisfiability lemma: for every standard conjunction $\phi^{\prime}$, if the satisfiability-reduction-condition of $\phi^{\prime}$ is met, then $\phi^{\prime}$ is satisfiable. Here, a standard conjunction $\phi^{\prime}$ is equivalent to the negation of a standard disjunction $\phi$, and the satisfiability-reduction-condition of $\phi^{\prime}$ is equivalent to the negation of the validity-reduction-condition of $\phi$.
To show the upward satisfiability lemma, the key is to find a pointed model satisfying a standard conjunction $\phi^{\prime}$, given its satisfiability-reduction-condition is met. To achieve this, we will use a method called the blueprint method.
3.2 Standard formulas and a normal form lemma
For every $n\in\mathbb{N}$, the set $\{-x\mid x\in\mathbb{N}\text{ and }1\leq x\leq n\}$ is called a negative indice set and the set $\{x\mid x\in\mathbb{N}\text{ and }0\leq x\leq n\}$ is called a positive indice set. Note negative indice sets are allowed to be empty, but positive indice sets are not.
We call a disjunction of propositional literals an elementary disjunction, and a conjunction of propositional literals an elementary conjunction.
Definition 8 (Standard disjunctions and conjunctions).
A standard disjunction is a formula in the form $\gamma\vee(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\phi_{j})$, where $\gamma$ is an elementary disjunction, $\mathbb{I}$ is a negative indice set, and $\mathbb{J}$ is a positive indice set such that (1) if $\mathbb{I}\neq\emptyset$, then $\langle[\mathrm{A}_{i}]\rangle\phi_{i}=\langle[\emptyset]\rangle\top$ for some $i\in\mathbb{I}$, and (2) $\langle[\mathrm{B}_{0}]\rangle\phi_{0}=\langle[\mathtt{AG}]\rangle\bot$.
A standard conjunction is a formula in the form $\gamma\wedge\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\wedge\bigwedge_{j\in\mathbb{J}}\neg\langle[\mathrm{B}_{j}]\rangle\phi_{j}$, where $\gamma$ is an elementary conjunction, $\mathbb{I}$ is a negative indice set, and $\mathbb{J}$ is a positive indice set such that (1) if $\mathbb{I}\neq\emptyset$, then $\langle[\mathrm{A}_{i}]\rangle\phi_{i}=\langle[\emptyset]\rangle\top$ for some $i\in\mathbb{I}$, and (2) $\neg\langle[\mathrm{B}_{0}]\rangle\phi_{0}=\neg\langle[\mathtt{AG}]\rangle\bot$.
It is easy to see that the negation of a standard disjunction is equivalent to a standard conjunction, and vice versa.
We define the modal depth of formulas as usual.
Lemma 1 (Normal form).
Let $\mathsf{X}\in\mathrm{ES}$. For every $\phi\in\Phi$, there is $\phi^{\prime}$ such that (1) $\vdash_{\mathsf{MCL}+\mathsf{X}}\phi\leftrightarrow\phi^{\prime}$, (2) $\phi$ and $\phi^{\prime}$ have the same modal depth, and (3) $\phi^{\prime}$ is in the form of $\chi_{0}\land\dots\land\chi_{k}$, where every $\chi_{i}$ is a standard disjunction.
[…]
The following notions will be used in stating the downward validity lemma and the upward derivability lemma.
Definition 9 (Sets of basic positive indices).
Let $\gamma\vee(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$ be a standard disjunction. $\mathbb{J}_{0}=\{j\in\mathbb{J}\mid\mathrm{B}_{i}=\mathtt{AG}\}$ is called the set of basic positive indices.
Definition 10 (Neat sets of negative indices).
Let $\mathsf{X}\in\mathrm{ES}$ and $\gamma\vee(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$ be a standard disjunction.
For every $\mathbb{I}^{\prime}\subseteq\mathbb{I}$, we say $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat if the following conditions are met:
  1. (1)
    for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $i\neq i^{\prime}$, then $\mathrm{A}_{i}\cap\mathrm{A}_{i^{\prime}}=\emptyset$;
  2. (2)
    if $\mathsf{S}\notin\mathsf{X}$, then $\mathbb{I}^{\prime}\neq\emptyset$;
  3. (3)
    if $\mathsf{I}\notin\mathsf{X}$, then for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $\mathrm{A}_{i}\neq\emptyset$ and $\mathrm{A}_{i^{\prime}}\neq\emptyset$, then $i=i^{\prime}$.
The following fact can be easily shown and we skip its proof:
Fact 2.
Let $\gamma\vee(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$ be a standard disjunction. Then, for all $\mathsf{X}\in\mathrm{ES}$ and $\mathbb{I}^{\prime},\mathbb{I}^{\prime\prime}\subseteq\mathbb{I}$, if $\mathbb{I}^{\prime}\subseteq\mathbb{I}^{\prime\prime}$ and $\mathbb{I}^{\prime\prime}$ is $\mathsf{X}$-neat, then $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat.
Roughly, the point of the fact is that neatness is conversely monotonic for negative indice sets.
3.3 Downward validity lemma
3.3.1 Blueprints and their realization
Definition 11 (Blueprints).
A blueprint is a tuple $\mathtt{BP}=(\mathtt{AG},\mathtt{AC},\mathtt{list}_{\mathtt{AG}})$, where:
  • $\mathtt{AG}$ is a finite set of agents;
  • $\mathtt{AC}$ is a nonempty set of actions;
  • $\mathtt{list}_{\mathtt{AG}}:\mathtt{JA}_{\mathtt{AG}}\to{\mathcal{P}(\Phi)}$ is a listing funtion for $\mathtt{AG}$.
For any $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, intuitively, $\mathtt{list}_{\mathtt{AG}}(\sigma_{\mathtt{AG}})$ lists some formulas which $\sigma_{\mathtt{AG}}$ enables.
Definition 12 (Performable joint actions).
Let $\mathtt{BP}=(\mathtt{AG},\mathtt{AC},\mathtt{list}_{\mathtt{AG}})$ be a blueprint.
For all $\mathrm{C}\subseteq\mathtt{AG}$, define the listing function $\mathtt{list}_{\mathrm{C}}$ for $\mathrm{C}$ as follows: for all $\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}$,
$\mathtt{list}_{\mathrm{C}}(\sigma_{\mathrm{C}})=\bigcup\{\mathtt{list}_{\mathtt{AG}}(\sigma_{\mathtt{AG}})\mid\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}\text{ and }\sigma_{\mathrm{C}}\subseteq\sigma_{\mathtt{AG}}\}.$
For all $\mathrm{C}\subseteq\mathtt{AG}$, define the set of performable joint actions $\mathtt{PJA}_{\mathrm{C}}$ of $\mathrm{C}$ as follows:
$\mathtt{PJA}_{\mathrm{C}}=\{\sigma_{\mathrm{C}}\in\mathtt{JA}_{\mathrm{C}}\mid\mathtt{list}_{\mathrm{C}}({\sigma_{\mathrm{C}}})\neq\emptyset\}.$
We use $\mathtt{PJA}$ to denote $\bigcup\{\mathtt{PJA}_{\mathrm{C}}\mid\mathrm{C}\subseteq\mathtt{AG}\}$.
Definition 13 (Regular blueprints).
Let $\mathtt{BP}=(\mathtt{AG},\mathtt{AC},\mathtt{list}_{\mathtt{AG}})$ be a blueprint and $\mathsf{X}\in\mathrm{ES}$.
We say $\mathtt{BP}$ is $\mathsf{X}$-regular if the following conditions are met:
  1. (1)
    for all $\mathrm{C}\subseteq\mathtt{AG}$, ${\sigma_{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$, and $\chi\in\mathtt{list}_{\mathrm{C}}({\sigma_{\mathrm{C}}})$, $\chi$ is $\mathsf{X}$-satisfiable;
  2. (2)
    if $\mathsf{S}\in\mathsf{X}$, then $\mathtt{BP}$ is serial, that is, $\mathtt{PJA}_{\mathrm{C}}\neq\emptyset$ for all $\mathrm{C}\subseteq\mathtt{AG}$;
  3. (3)
    if $\mathsf{I}\in\mathsf{X}$, then $\mathtt{BP}$ is independent, that is, ${\sigma_{\mathrm{C}}}\cup{\sigma_{\mathrm{D}}}\in\mathtt{PJA}_{\mathrm{C}\cup\mathrm{D}}$ for all $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, ${\sigma_{\mathrm{C}}}\in\mathtt{PJA}_{\mathrm{C}}$, and ${\sigma_{\mathrm{D}}}\in\mathtt{PJA}_{\mathrm{D}}$;
  4. (4)
    if $\mathsf{D}\in\mathsf{X}$, then $\mathtt{BP}$ is deterministic, that is, $\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})$ is a singleton for all ${\sigma_{\mathtt{AG}}}\in\mathtt{PJA}_{\mathtt{AG}}$.
Definition 14 (Ensuring functions and enabling functions).
Let $\mathsf{X}\in\mathrm{ES}$ and $\mathtt{BP}=(\mathtt{AG},\mathtt{AC},\mathtt{list}_{\mathtt{AG}})$ be an $\mathsf{X}$-regular blueprint.
Define the ensuring function $\mathcal{A}_{\mathsf{X}}$ as follows: for all $\mathrm{C}\subseteq\mathtt{AG}$ and ${\sigma_{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$,
${\mathcal{A}_{\mathsf{X}}({\sigma_{\mathrm{C}}})}=\{\psi\in\Phi\mid\text{ for all }\phi\in\mathtt{list}_{\mathrm{C}}({\sigma_{\mathrm{C}}}),\phi\vDash_{\mathsf{MCL}+\mathsf{X}}\psi\}.$
Define the enabling function $\mathcal{E}_{\mathsf{X}}$ as follows: for all $\mathrm{C}\subseteq\mathtt{AG}$ and ${\sigma_{\mathrm{C}}}\in\mathtt{PJA}_{\mathrm{C}}$,
${\mathcal{E}_{\mathsf{X}}({\sigma_{\mathrm{C}}})}=\{\psi\in\Phi\mid\text{ for some }\phi\in\mathtt{list}_{\mathrm{C}}({\sigma_{\mathrm{C}}}),\phi\vDash_{\mathsf{MCL}+\mathsf{X}}\psi\}.$
Intuitively, ${\sigma_{\mathrm{C}}}$ ensures the formulas in ${\mathcal{A}_{\mathsf{X}}({\sigma_{\mathrm{C}}})}$, and enables the formulas in ${\mathcal{E}_{\mathsf{X}}({\sigma_{\mathrm{C}}})}$.
Definition 15 (Realization of blueprints).
Let $\mathsf{X}\in\mathrm{ES}$, $\mathtt{BP}=(\mathtt{AG},\mathtt{AC}_{0},\mathtt{list}_{\mathtt{AG}})$ be an $\mathsf{X}$-regular blueprint, and $\gamma$ be an $\mathsf{X}$-satisfiable elementary conjunction.
We say a pointed $\mathsf{X}$-model $(\mathtt{M},s_{0})$, where $\mathtt{M}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathtt{AP},\mathsf{label})$, realizes $\mathtt{BP}$ and $\gamma$ if the following conditions are met:
  1. (1)
    $\mathtt{AC}_{0}\subseteq\mathtt{AC}$;
  2. (2)
    $\mathtt{aja}_{\mathrm{C}}(s_{0})=\mathtt{PJA}_{\mathrm{C}}$ for every $\mathrm{C}\subseteq\mathtt{AG}$;
  3. (3)
    $\mathtt{M},s_{0}\Vdash\gamma$;
  4. (4)
    ${\sigma_{\mathrm{C}}}\vartriangleright_{(\mathtt{M},s_{0})}\psi$ for every $\mathrm{C}\subseteq\mathtt{AG}$, ${\sigma_{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$ and $\psi\in{\mathcal{A}_{\mathsf{X}}({\sigma_{\mathrm{C}}})}$;
  5. (5)
    ${\sigma_{\mathrm{C}}}\ltimes_{(\mathtt{M},s_{0})}\psi$ for every $\mathrm{C}\subseteq\mathtt{AG}$, ${\sigma_{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$ and $\psi\in{\mathcal{E}_{\mathsf{X}}({\sigma_{\mathrm{C}}})}$.
Theorem 1 (Realizability of blueprints).
Let $\mathsf{X}\in\mathrm{ES}$. For all $\mathsf{X}$-regular blueprint $\mathtt{BP}$ and $\mathsf{X}$-satisfiable elementary conjunction $\gamma$, there is a pointed $\mathsf{X}$-model $(\mathtt{M},s_{0})$ realizing them.
Proof.
Let $\mathtt{BP}=(\mathtt{AG},\mathtt{AC}_{0},\mathtt{list}_{\mathtt{AG}})$ be an $\mathsf{X}$-regular blueprint, and $\gamma$ be an $\mathsf{X}$-satisfiable elementary conjunction.
Let $\{(\mathtt{M}^{\lambda}_{\phi},s^{\lambda}_{\phi})\mid\lambda\in\mathtt{PJA}_{\mathtt{AG}}\text{ and }\phi\in\mathtt{list}_{\mathtt{AG}}(\lambda)\}$, where $\mathtt{M}^{\lambda}_{\phi}=(\mathtt{ST}^{\lambda}_{\phi},\mathtt{AG},\mathtt{AC}^{\lambda}_{\phi},{\mathsf{out}_{\mathtt{AG}}}^{\lambda}_{\phi},\mathtt{AP},\mathsf{label}^{\lambda}_{\phi})$, be a set of pointed $\mathsf{X}$-models meeting the following conditions:
  1. (1)
    for every $\lambda\in\mathtt{PJA}_{\mathtt{AG}}$ and $\phi\in\mathtt{list}_{\mathtt{AG}}(\lambda)$, $\mathtt{M}^{\lambda}_{\phi},s^{\lambda}_{\phi}\Vdash\phi$;
  2. (2)
    all $\mathtt{ST}^{\lambda}_{\phi}$ are pairwise disjoint;
  3. (3)
    all $\mathtt{AC}^{\lambda}_{\phi}$ and $\mathtt{AC}_{0}$ are pairwise disjoint.
Note $\{(\mathtt{M}^{\lambda}_{\phi},s^{\lambda}_{\phi})\mid\lambda\in\mathtt{PJA}_{\mathtt{AG}}\text{ and }\phi\in\mathtt{list}_{\mathtt{AG}}(\lambda)\}$ is empty if $\mathtt{PJA}_{\mathtt{AG}}$ is empty.
Let $s_{0}$ be a state not in any of these models.
Define a pointed model $(\mathtt{M},s_{0})$, where $\mathtt{M}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathtt{AP},\mathsf{label})$, as follows:
  • $\mathtt{ST}=\{s_{0}\}\cup\bigcup\{\mathtt{ST}^{\lambda}_{\phi}\mid\lambda\in\mathtt{PJA}_{\mathtt{AG}}\text{ and }\phi\in\mathtt{list}_{\mathtt{AG}}(\lambda)\}$;
  • $\mathtt{AC}=\mathtt{AC}_{0}\cup\bigcup\{\mathtt{AC}^{\lambda}_{\phi}\mid\lambda\in\mathtt{PJA}_{\mathtt{AG}}\text{ and }\phi\in\mathtt{list}_{\mathtt{AG}}(\lambda)\}$;
  • $\mathsf{out}_{\mathtt{AG}}:\mathtt{ST}\times\mathtt{JA}_{\mathtt{AG}}\to{\mathcal{P}(\mathtt{ST})}$ such that for all $s\in\mathtt{ST}$ and ${\sigma_{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}$:
    ${\mathsf{out}_{\mathtt{AG}}(s,{\sigma_{\mathtt{AG}}})}=\begin{cases}{{\mathsf{out}_{\mathtt{AG}}}^{\lambda}_{\phi}(s,{\sigma_{\mathtt{AG}}}})&\text{if }s\in{\mathtt{ST}^{\lambda}_{\phi}}\text{ and }{\sigma_{\mathtt{AG}}}\in{\mathtt{JA}_{\mathtt{AG}}}^{\lambda}_{\phi}\\ &\text{for some }\lambda\in\mathtt{PJA}_{\mathtt{AG}}\text{ and }\phi\in\mathtt{list}_{\mathtt{AG}}(\lambda)\\ \{s^{{\sigma_{\mathtt{AG}}}}_{\phi}\mid\phi\in\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})\}&\text{if }s=s_{0}\text{ and }{\sigma_{\mathtt{AG}}}\in\mathtt{PJA}_{\mathtt{AG}}\\ \emptyset&\text{otherwise}\end{cases}$
    where ${\mathtt{JA}_{\mathtt{AG}}}^{\lambda}_{\phi}=\{\delta:\mathtt{AG}\to\mathtt{AC}^{\lambda}_{\phi}\}$;
  • $\mathsf{label}:\mathtt{ST}\to{\mathcal{P}(\mathtt{AP})}$ such that for all $s\in\mathtt{ST}$:
    ${\mathsf{label}(s)}=\begin{cases}{\mathsf{label}^{\lambda}_{\phi}(s)}&\text{if}s\in{\mathtt{ST}^{\lambda}_{\phi}}\text{ for some }\lambda\in\mathtt{PJA}_{\mathtt{AG}}\text{ and }\phi\in\mathtt{list}_{\mathtt{AG}}(\lambda)\\ \{p\mid\text{$p$ is a conjunct of $\gamma$}\}&\text{if }s=s_{0}\\ \end{cases}$
We claim $(\mathtt{M},s_{0})$ is a pointed $\mathsf{X}$-model realizing $\mathtt{BP}$ and $\gamma$.
It is easy to verify the first two conditions given in the definition of realization of blueprints:
  1. (1)
    $\mathtt{AC}_{0}\subseteq\mathtt{AC}$;
  2. (2)
    $\mathtt{aja}_{\mathrm{C}}(s_{0})=\mathtt{PJA}_{\mathrm{C}}$ for every $\mathrm{C}\subseteq\mathtt{AG}$.
Then, it is easy to check that $(\mathtt{M},s_{0})$ is a pointed $\mathsf{X}$-model.
Let $\lambda\in\mathtt{PJA}_{\mathtt{AG}}$ and $\phi\in\mathtt{list}_{\mathtt{AG}}(\lambda)$. It can be easily shown the generated sub-model of $\mathtt{M}^{\lambda}_{\phi}$ at $s^{\lambda}_{\phi}$ is also the generated sub-model of $\mathtt{M}$ at $s^{\lambda}_{\phi}$. Then for every $\psi\in\Phi$: $\mathtt{M}^{\lambda}_{\phi},s^{\lambda}_{\phi}\Vdash\psi$ if and only if $\mathtt{M},s^{\lambda}_{\phi}\Vdash\psi$.
Now we show the last three conditions given in the definition of realization of blueprints hold.
  1. (3)
    It is easy to see $\mathtt{M},s_{0}\Vdash\gamma$.
  2. (4)
    Let $\mathrm{C}\subseteq\mathtt{AG}$, ${\sigma_{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$ and $\psi\in{\mathcal{A}_{X}({\sigma_{\mathrm{C}}})}$. We want to show ${\sigma_{\mathrm{C}}}\vartriangleright_{(\mathtt{M},s_{0})}\psi$.
    Then, for all $\phi\in\mathtt{list}_{\mathrm{C}}({\sigma_{\mathrm{C}}})$, $\phi\vDash_{\mathsf{MCL}+\mathsf{X}}\psi$. Then for all ${\sigma_{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}$ and $\phi\in\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})$, if ${\sigma_{\mathrm{C}}}\subseteq{\sigma_{\mathtt{AG}}}$, then $\phi\vDash_{{\mathsf{MCL}+\mathsf{X}}}\psi$.
    Let $s\in\mathsf{out}_{\mathrm{C}}(s_{0},\sigma_{\mathrm{C}})$. Then $s\in\mathsf{out}_{\mathtt{AG}}(s_{0},\sigma_{\mathtt{AG}})$ for some $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathtt{AG}}\subseteq\sigma_{\mathrm{C}}$. Then, $s=s^{\sigma_{\mathtt{AG}}}_{\phi}$ for some $\phi\in\mathtt{list}_{\mathtt{AG}}(\sigma_{\mathtt{AG}})$. Note $\mathtt{M}^{\sigma_{\mathtt{AG}}}_{\phi},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\phi$. Then $\mathtt{M},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\phi$. Then $\mathtt{M},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\psi$, that is, $\mathtt{M},s\Vdash\psi$. Therefore, ${\sigma_{\mathrm{C}}}\vartriangleright_{(\mathtt{M},s_{0})}\psi$.
  3. (5)
    Let $\mathrm{C}\subseteq\mathtt{AG}$, ${\sigma_{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$ and $\psi\in{\mathcal{E}_{\mathsf{X}}({\sigma_{\mathrm{C}}})}$. We want to show ${\sigma_{\mathrm{C}}}\ltimes_{(\mathtt{M},s_{0})}\psi$.
    Then, for some $\phi\in\mathtt{list}_{\mathrm{C}}({\sigma_{\mathrm{C}}})$, $\phi\vDash_{\mathsf{MCL}+\mathsf{X}}\psi$. Then for some ${\sigma_{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}$, ${\sigma_{\mathrm{C}}}\subseteq{\sigma_{\mathtt{AG}}}$ and $\phi\in\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})$. Then ${\sigma_{\mathtt{AG}}}\in\mathtt{PJA}_{\mathtt{AG}}$. Then the pointed $\mathsf{X}$-model $(\mathtt{M}^{\sigma_{\mathtt{AG}}}_{\phi},s^{\sigma_{\mathtt{AG}}}_{\phi})$ is defined and $\mathtt{M}^{\sigma_{\mathtt{AG}}}_{\phi},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\phi$. Then $\mathtt{M},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\phi$. Then, $\mathtt{M},s^{\sigma_{\mathtt{AG}}}_{\phi}\Vdash\psi$. Note $s^{\sigma_{\mathtt{AG}}}_{\phi}\in\mathsf{out}_{\mathtt{AG}}(s_{0},\sigma_{\mathtt{AG}})$. Then $s^{\sigma_{\mathtt{AG}}}_{\phi}\in\mathsf{out}_{\mathrm{C}}(s_{0},\sigma_{\mathrm{C}})$. Therefore, ${\sigma_{\mathrm{C}}}\ltimes_{(\mathtt{M},s_{0})}\psi$.
3.3.2 The lemma
Let $\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$ be a standard disjunction. We use $\mathbb{J}_{0}$ to indicate $\{j\in\mathbb{J}\mid\mathrm{B}_{j}=\mathtt{AG}\}$.
Lemma 2 (Downward validity).
Let $\mathsf{X}\in\mathrm{ES}$ and $\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$ be a standard disjunction. Then, if $\models_{{\mathsf{MCL}+\mathsf{X}}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$, then the following condition, called the $\mathsf{X}$-validity-reduction-condition, is met:
  1. (1)
    $\models_{{\mathsf{MCL}+\mathsf{X}}}\gamma$, or
  2. (2)
    the following two conditions hold:
    1. (a)
      if $\mathsf{D}\notin\mathsf{X}$, then there is $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$ such that $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat, $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, and $\models_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\to\psi_{j}$;
    2. (b)
      if $\mathsf{D}\in\mathsf{X}$, then there is $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$ such that $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat, $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, and $\models_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\to(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$.
Proof.
Assume the $\mathsf{X}$-validity-reduction-condition is not met. Then, it can be verified that the following condition, called the $\mathsf{X}$-satisfiability-reduction-condition, is met:
  1. (1)
    $\neg\gamma$ is $\mathsf{X}$-satisfiable, and
  2. (2)
    the following two conditions hold:
    1. (a)
      if $\mathsf{D}\notin\mathsf{X}$, then for all $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$, if $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat and $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, then $\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\wedge\neg\psi_{j}$ is $\mathsf{X}$-satisfiable;
    2. (b)
      if $\mathsf{D}\in\mathsf{X}$, then for all $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$, if $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat and $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, then $\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\wedge\neg\psi_{j}\wedge\bigwedge_{k\in\mathbb{J}_{0}}\neg\psi_{k}$ is $\mathsf{X}$-satisfiable.
Let $\gamma^{\prime}$ be an $\mathsf{X}$-satisfiable elementary conjunction, equivalent to $\neg\gamma$. It suffices to show the standard conjunction $\gamma^{\prime}\wedge\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\wedge\bigwedge_{j\in\mathbb{J}}\neg\langle[\mathrm{B}_{j}]\rangle\psi_{j}$ is $\mathsf{X}$-satisfiable.
In the sequel, first, we define a blueprint, from which we can get a pointed $\mathsf{X}$-model; second, we show the pointed $\mathsf{X}$-model satisfies $\gamma^{\prime}\wedge\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\wedge\bigwedge_{j\in\mathbb{J}}\neg\langle[\mathrm{B}_{j}]\rangle\psi_{j}$.
Define two functions $\mathtt{support}:\mathtt{JA}\to{\mathcal{P}(\mathbb{I})}$ and $\mathtt{impeach}:\mathtt{JA}\to\mathbb{J}$ as follows:
  • for all $\mathrm{C}\subseteq\mathtt{AG}$ and ${\sigma_{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$:
    $\mathtt{support}({\sigma_{\mathrm{C}}})=\{i\in\mathbb{I}\mid\mathrm{A}_{i}\subseteq\mathrm{C}\text{ and }{{\sigma_{\mathrm{C}}}(a)}=i\text{ for all }a\in\mathrm{A}_{i}\}.$
  • for all $\mathrm{C}\subseteq\mathtt{AG}$ and ${\sigma_{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$:
    ${\mathtt{impeach}({\sigma_{\mathrm{C}}})}=\Big{(}\sum\{{{\sigma_{\mathrm{C}}}(a)}\in\mathbb{J}\mid a\in\mathrm{C}\}\Big{)}\bmod n,$
    where $\sum$ is the sum operation and $n$ is the cardinality of $\mathbb{J}$.
What follows are some facts about the function $\mathtt{support}$:
  1. (1)
    Claim 1: for all $\mathrm{C}\subseteq\mathtt{AG}$ and ${\sigma_{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$, $\bigcup_{i\in\mathtt{support}({\sigma_{\mathrm{C}}})}\mathrm{A}_{i}\subseteq\mathrm{C}$, and ${\mathtt{support}({\sigma_{\mathrm{C}}})}$ is both $\mathsf{S}\mathsf{I}$-neat and $\mathsf{S}\mathsf{I}\mathsf{D}$-neat.
    Let $\mathrm{C}\subseteq\mathtt{AG}$ and ${\sigma_{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$. From the definition of $\mathtt{support}$, we can see $\bigcup_{i\in\mathtt{support}({\sigma_{\mathrm{C}}})}\mathrm{A}_{i}\subseteq\mathrm{C}$. Let $i,i^{\prime}\in\mathtt{support}({\sigma_{\mathrm{C}}})$ such that $i\neq i^{\prime}$. To show ${\mathtt{support}({\sigma_{\mathrm{C}}})}$ is both $\mathsf{S}\mathsf{I}$-neat and $\mathsf{S}\mathsf{I}\mathsf{D}$-neat, it suffices to show $\mathrm{A}_{i}\cap\mathrm{A}_{i^{\prime}}=\emptyset$. Assume $\mathrm{A}_{i}\cap\mathrm{A}_{i^{\prime}}\neq\emptyset$. Let $a\in\mathrm{A}_{i}\cap\mathrm{A}_{i^{\prime}}$. By the definition of $\mathtt{support}$, $\sigma_{\mathrm{C}}(a)=i$ and ${{\sigma_{\mathrm{C}}}(a)}=i^{\prime}$. Then $i=i^{\prime}$. We have a contradiction.
  2. (2)
    Claim 2: for all $\mathrm{C},\mathrm{C}^{\prime}\subseteq\mathtt{AG}$, ${\sigma_{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$, and ${\sigma_{\mathrm{C}^{\prime}}}\in\mathtt{JA}_{\mathrm{C}}$, if $\mathrm{C}\subseteq\mathrm{C}^{\prime}$ and ${\sigma_{\mathrm{C}}}\subseteq{\sigma_{\mathrm{C}^{\prime}}}$, then ${\mathtt{support}({\sigma_{\mathrm{C}}})}\subseteq\mathtt{support}({\sigma_{\mathrm{C}^{\prime}}})$.
    Let $\mathrm{C},\mathrm{C}^{\prime}\subseteq\mathtt{AG}$, ${\sigma_{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$, and ${\sigma_{\mathrm{C}^{\prime}}}\in\mathtt{JA}_{\mathrm{C}}$ such that $\mathrm{C}\subseteq\mathrm{C}^{\prime}$ and ${\sigma_{\mathrm{C}}}\subseteq{\sigma_{\mathrm{C}^{\prime}}}$. Let $i\in\mathtt{support}({\sigma_{\mathrm{C}}})$. Then $\mathrm{A}_{i}\subseteq\mathrm{C}$ and ${{\sigma_{\mathrm{C}}}(a)}=i$ for all $a\in\mathrm{A}_{i}$. Then $\mathrm{A}_{i}\subseteq\mathrm{C}^{\prime}$ and ${{\sigma_{\mathrm{C}^{\prime}}}(a)}=i$ for all $a\in\mathrm{A}_{i}$. Then, $i\in\mathtt{support}({\sigma_{\mathrm{C}^{\prime}}})$.
  3. (3)
    Claim 3: if $\mathbb{I}\neq\emptyset$, then for all $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, $\mathtt{support}({\sigma_{\mathtt{AG}}})\neq\emptyset$.
    Assume $\mathbb{I}\neq\emptyset$. By the definition of standard disjunctions, there is $i\in\mathbb{I}$ such that $\mathrm{A}_{i}=\emptyset$. Then $i\in\mathtt{support}({\sigma_{\mathtt{AG}}})$.
Define a blueprint $\mathtt{BP}=(\mathtt{AG},\mathtt{AC}_{0},\mathtt{list}_{\mathtt{AG}})$ as follows:
  • $\mathtt{AC}=\mathbb{I}\cup\mathbb{J}$;
  • if $\mathsf{D}\notin\mathsf{X}$, then for every ${\sigma_{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}$:
    $\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})=$
    $\begin{cases}\Big{\{}\bigwedge_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\phi_{i}\wedge\neg\psi_{j}\mid j\in\mathbb{J}\text{ and }\bigcup_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}\Big{\}}&\parbox[t]{120.00018pt}{ \text{if $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat} }\\ \emptyset&\text{otherwise}\end{cases}$
  • if $\mathsf{D}\in\mathsf{X}$, then for every ${\sigma_{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}$:
    $\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})=$
    $\begin{cases}\Big{\{}\bigwedge_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\phi_{i}\wedge\neg\psi_{\mathtt{impeach}({\sigma_{\mathtt{AG}}})}\wedge\bigwedge_{k\in\mathbb{J}_{0}}\neg\psi_{k}\Big{\}}&\parbox[t]{200.0003pt}{ \text{if $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat and} }\\ &\parbox[t]{200.0003pt}{ \text{$\bigcup_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\mathrm{A}_{i}\subseteq\mathrm{B}_{\mathtt{impeach}({\sigma_{\mathtt{AG}}})}$} }\vspace{15pt}\\ \Big{\{}\bigwedge_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\phi_{i}\wedge\neg\psi_{0}\wedge\bigwedge_{k\in\mathbb{J}_{0}}\neg\psi_{k}\Big{\}}&\parbox[t]{200.0003pt}{ \text{if $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat and} }\\ &\parbox[t]{200.0003pt}{ \text{$\bigcup_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\mathrm{A}_{i}\not\subseteq\mathrm{B}_{\mathtt{impeach}({\sigma_{\mathtt{AG}}})}$} }\vspace{15pt}\\ \emptyset&\text{otherwise}\end{cases}$
The following fact can be easily seen:
  1. (4)
    Claim 4: For every $\sigma_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$, $\mathtt{list}_{\mathtt{AG}}(\sigma_{\mathtt{AG}})\neq\emptyset$ if and only if $\mathtt{support}(\sigma_{\mathtt{AG}})$ is $\mathsf{X}$-neat.
We claim $\mathtt{BP}$ is $\mathsf{X}$-regular.
  1. (1)
    For all ${\sigma_{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}$ and $\chi\in\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})$, $\chi$ is $\mathsf{X}$-satisfiable.
    Let ${\sigma_{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}$ and $\chi\in\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})$.
    • Assume $\mathsf{D}\notin\mathsf{X}$. Then there is $j\in\mathbb{J}$ such that $\chi=\bigwedge_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\phi_{i}\wedge\neg\psi_{j}$, where $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat and $\bigcup_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$. By the $\mathsf{X}$-satisfiability-reduction-condition, $\chi$ is $\mathsf{X}$-satisfiable.
    • Assume $\mathsf{D}\in\mathsf{X}$. Then either (a) $\chi=\bigwedge_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\phi_{i}\wedge\neg\psi_{\mathtt{impeach}({\sigma_{\mathtt{AG}}})}\wedge\bigwedge_{k\in\mathbb{J}_{0}}\neg\psi_{k}$, where $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat and $\bigcup_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\mathrm{A}_{i}\subseteq\mathrm{B}_{\mathtt{impeach}({\sigma_{\mathtt{AG}}})}$, or (b) $\chi=\bigwedge_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\phi_{i}\wedge\neg\psi_{0}\wedge\bigwedge_{k\in\mathbb{J}_{0}}\neg\psi_{k}$, where $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat and $\bigcup_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}$ $\mathrm{A}_{i}\not\subseteq\mathrm{B}_{\mathtt{impeach}({\sigma_{\mathtt{AG}}})}$. Assume (a). By the $\mathsf{X}$-satisfiablity-redution-condition, $\chi$ is $\mathsf{X}$-satisfiable. Assume (b). Note $\mathrm{B}_{0}=\mathtt{AG}$. Then $\bigcup_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\mathrm{A}_{i}\subseteq\mathrm{B}_{0}$. By the $\mathsf{X}$-satisfiablity-redution-condition, $\chi$ is $\mathsf{X}$-satisfiable.
  2. (2)
    If $\mathsf{S}\in\mathsf{X}$, then $\mathtt{BP}$ is serial.
    Assume $\mathsf{S}\in\mathsf{X}$. Let $\mathrm{C}\subseteq\mathtt{AG}$. It suffices to show $\mathtt{PJA}_{\mathrm{C}}\neq\emptyset$. Let ${\sigma_{\mathrm{C}}}\in\mathtt{JA}_{\mathrm{C}}$ such that ${{\sigma_{\mathrm{C}}}(a)}=0$ for all $a\in\mathrm{C}$. It suffices to show $\mathtt{list}_{\mathrm{C}}(\sigma_{\mathrm{C}})\neq\emptyset$. Let ${\sigma_{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}$ such that ${{\sigma_{\mathtt{AG}}}(a)}=0$ for all $a\in\mathtt{AG}$. Then ${\sigma_{\mathrm{C}}}\subseteq{\sigma_{\mathtt{AG}}}$. It suffices to show $\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})\neq\emptyset$. It can be checked $\mathtt{support}({\sigma_{\mathtt{AG}}})=\{i\in\mathbb{I}\mid\mathrm{A}_{i}=\emptyset\}$ and $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat. By Claim (4), $\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})\neq\emptyset$.
  3. (3)
    If $\mathsf{I}\in\mathsf{X}$, then $\mathtt{BP}$ is independent.
    Assume $\mathsf{I}\in\mathsf{X}$. Let $\mathrm{C},\mathrm{D}\subseteq\mathtt{AG}$ such that $\mathrm{C}\cap\mathrm{D}=\emptyset$, ${\sigma_{\mathrm{C}}}\in\mathtt{PJA}_{\mathrm{C}}$, and ${\sigma_{\mathrm{D}}}\in\mathtt{PJA}_{\mathrm{D}}$. We want to show ${\sigma_{\mathrm{C}}}\cup{\sigma_{\mathrm{D}}}\in\mathtt{PJA}_{\mathrm{C}\cup\mathrm{D}}$. It suffices to show $\mathtt{list}_{\mathrm{C}\cup\mathrm{D}}(\sigma_{\mathrm{C}}\cup\sigma_{\mathrm{D}})\neq\emptyset$. Let ${\sigma_{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}$ such that ${\sigma_{\mathrm{C}}}\cup{\sigma_{\mathrm{D}}}\subseteq{\sigma_{\mathtt{AG}}}$ and ${{\sigma_{\mathtt{AG}}}(a)}=0$ for all $a\in\mathtt{AG}-(\mathrm{C}\cup\mathrm{D})$. It suffices to show $\mathtt{list}_{\mathtt{AG}}(\sigma_{\mathtt{AG}})\neq\emptyset$.
    By Claim (1), $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is both $\mathsf{S}\mathsf{I}$-neat and $\mathsf{S}\mathsf{I}\mathsf{D}$-neat.
    • Assume $\mathsf{S}\in\mathsf{X}$. Then, it can be checked $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat. By Claim (4), $\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})\neq\emptyset$.
    • Assume $\mathsf{S}\notin\mathsf{X}$.
      • Assume $\mathtt{support}({\sigma_{\mathtt{AG}}})\neq\emptyset$. It can be verified $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat. By Claim (4), $\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})\neq\emptyset$.
      • Assume $\mathtt{support}({\sigma_{\mathtt{AG}}})=\emptyset$. We want to show this case is impossible. By Claim (3), $\mathbb{I}=\emptyset$. Therefore, $\mathtt{support}(\lambda_{\mathtt{AG}})=\emptyset$ for all $\lambda_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$. Note ${\sigma_{\mathrm{C}}}\in\mathtt{PJA}_{\mathrm{C}}$. Then $\mathtt{list}_{\mathrm{C}}(\sigma_{\mathrm{C}})\neq\emptyset$. Then there is $\lambda_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that $\sigma_{\mathrm{C}}\subseteq\lambda_{\mathtt{AG}}$ and $\mathtt{list}_{\mathtt{AG}}(\lambda_{\mathtt{AG}})\neq\emptyset$. By Claim (4), $\mathtt{support}(\lambda_{\mathtt{AG}})$ is $\mathsf{X}$-neat. Note $\mathsf{S}\notin\mathsf{X}$. Then $\mathtt{support}(\lambda_{\mathtt{AG}})\neq\emptyset$. We have a contradiction.
  4. (4)
    If $\mathsf{D}\in\mathsf{X}$, then $\mathtt{BP}$ is deterministic.
    Assume $\mathsf{D}\in\mathsf{X}$. From the definition of $\mathtt{list}_{\mathtt{AG}}$, we can see that for every $\sigma_{\mathtt{AG}}\in\mathtt{PJA}_{\mathtt{AG}}$, $\mathtt{list}_{\mathtt{AG}}(\sigma_{\mathtt{AG}})$ is a singleton. Then, $\mathtt{BP}$ is deterministic.
By Theorem 1, there is a pointed $\mathsf{X}$-model $(\mathtt{M},s)$, where $\mathtt{M}=(\mathtt{ST},\mathtt{AG},\mathtt{AC},\mathsf{out}_{\mathtt{AG}},\mathtt{AP},\mathsf{label})$, realizing $\mathtt{BP}$ and $\gamma^{\prime}$. We claim $\mathtt{M},s\Vdash\gamma^{\prime}\land\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\wedge\bigwedge_{j\in\mathbb{J}}\neg\langle[\mathrm{B}_{j}]\rangle\psi_{j}$.
Let $i\in\mathbb{I}$. We want to show $\mathtt{M},s\Vdash\langle[\mathrm{A}_{i}]\rangle\phi_{i}$. Let ${\sigma_{\mathrm{A}_{i}}}\in\mathtt{JA}_{\mathrm{A}_{i}}$ such that ${{\sigma_{\mathrm{A}_{i}}}(a)}=i$ for all $a\in\mathrm{A}_{i}$. It suffices to show $\sigma_{\mathrm{A}_{i}}\in\mathtt{PJA}_{\mathrm{A}_{i}}$ and ${\sigma_{\mathrm{A}_{i}}}\vartriangleright_{(\mathtt{M},s)}\phi_{i}$.
Let ${\sigma_{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}$ such that ${{\sigma_{\mathtt{AG}}}(a)}=i$ for all $a\in\mathtt{AG}$. Then ${\sigma_{\mathrm{A}_{i}}}\subseteq{\sigma_{\mathtt{AG}}}$. It can be checked $\mathtt{support}({\sigma_{\mathtt{AG}}})=\{i^{\prime}\in\mathbb{I}\mid\mathrm{A}_{i^{\prime}}=\emptyset\text{ or }i^{\prime}=i\}$ and $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat. By Claim (4), $\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})\neq\emptyset$. Then $\mathtt{list}_{\mathrm{A}_{i}}(\sigma_{\mathrm{A}_{i}})\neq\emptyset$. Then $\sigma_{\mathrm{A}_{i}}\in\mathtt{PJA}_{\mathrm{A}_{i}}$.
We claim $\phi_{i}\in{\mathcal{A}_{\mathsf{X}}({\sigma_{\mathrm{A}_{i}}})}$. Let $\chi\in\mathtt{list}_{\mathrm{A}_{i}}(\sigma_{\mathrm{A}_{i}})$. Then $\chi\in\mathtt{list}_{\mathtt{AG}}(\lambda_{\mathtt{AG}})$ for some $\lambda_{\mathtt{AG}}\in\mathtt{JA}_{\mathtt{AG}}$ such that ${\sigma_{\mathrm{A}_{i}}}\subseteq\lambda_{\mathtt{AG}}$. Note $i\in\mathtt{support}({\sigma_{\mathrm{A}_{i}}})$. By Claim (2), $i\in\mathtt{support}(\lambda_{\mathtt{AG}})$. From the definition of $\mathtt{list}_{\mathtt{AG}}$, we can see $\phi_{i}$ is a conjunct of $\chi$. Then $\chi\vDash_{\mathsf{MCL}+\mathsf{X}}\phi_{i}$. Then $\phi_{i}\in{\mathcal{A}_{\mathsf{X}}({\sigma_{\mathrm{A}_{i}}})}$.
By theorem 1, ${\sigma_{\mathrm{A}_{i}}}\vartriangleright_{(\mathtt{M},s)}\phi_{i}$.
Let $j\in\mathbb{J}$. We want to show $\mathtt{M},s\Vdash\neg\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Let ${\sigma_{\mathrm{B}_{j}}}\in\mathtt{PJA}_{\mathrm{B}_{j}}$. It suffices to show ${\sigma_{\mathrm{B}_{j}}}\ltimes_{(\mathtt{M},s)}\neg\psi_{j}$.
Then, $\mathtt{list}_{\mathrm{B}_{i}}({\sigma_{\mathrm{B}_{j}}})\neq\emptyset$. Then there is $\lambda_{\mathtt{AG}}\in\mathtt{PJA}_{\mathtt{AG}}$ such that ${\sigma_{\mathrm{B}_{j}}}\subseteq\lambda_{\mathtt{AG}}$ and $\mathtt{list}_{\mathtt{AG}}(\lambda_{\mathtt{AG}})\neq\emptyset$. By Claim (4), $\mathtt{support}(\lambda_{\mathtt{AG}})$ is $\mathsf{X}$-neat. By Claim (2) and Fact 2, $\mathtt{support}({\sigma_{\mathrm{B}_{j}}})$ is $\mathsf{X}$-neat.
  • Assume $\mathsf{D}\notin\mathsf{X}$. Let ${\sigma_{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}$ such that ${\sigma_{\mathrm{B}_{j}}}\subseteq{\sigma_{\mathtt{AG}}}$ and ${{\sigma_{\mathtt{AG}}}(\overline{b})}=j$ for all $\overline{b}\in\mathtt{AG}-\mathrm{B}_{j}$.
    We claim $\mathtt{support}({\sigma_{\mathtt{AG}}})=\mathtt{support}({\sigma_{\mathrm{B}_{j}}})$. By Claim (2), $\mathtt{support}({\sigma_{\mathrm{B}_{j}}})\subseteq\mathtt{support}({\sigma_{\mathtt{AG}}})$. Let $i\in\mathtt{support}({\sigma_{\mathtt{AG}}})$. Then ${{\sigma_{\mathtt{AG}}}(a)}=i$ for all $a\in\mathrm{A}_{i}$.
    • Assume $\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$. Note ${\sigma_{\mathrm{B}_{j}}}\subseteq{\sigma_{\mathtt{AG}}}$. Then ${{\sigma_{\mathrm{B}_{j}}}(a)}=i$ for all $a\in\mathrm{A}_{i}$. Then $i\in\mathtt{support}({\sigma_{\mathrm{B}_{j}}})$.
    • Assume $\mathrm{A}_{i}\nsubseteq\mathrm{B}_{j}$. We want to show this case is impossible. Let $\overline{b}\in\mathrm{A}_{i}-\mathrm{B}_{j}$. Then $\overline{b}\in\mathtt{AG}-\mathrm{B}_{j}$. Then ${{\sigma_{\mathtt{AG}}}(\overline{b})}=j$. Note $\overline{b}\in\mathrm{A}_{i}$. Then ${{\sigma_{\mathtt{AG}}}(\overline{b})}=i$. Then $j=i$. Note $i\in\mathbb{I}$ and $j\in\mathbb{J}$. Then $\mathbb{I}\cap\mathbb{J}\neq\emptyset$. We have a contradiction.
    Then $\mathtt{support}({\sigma_{\mathtt{AG}}})=\mathtt{support}({\sigma_{\mathrm{B}_{j}}})$.
    Note $\mathtt{support}({\sigma_{\mathrm{B}_{j}}})$ is $\mathsf{X}$-neat. Then $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat. By Claim (1), $\bigcup_{i\in\mathtt{support}(\mathrm{B}_{j})}\mathrm{A}_{i}$ $\subseteq\mathrm{B}_{j}$. Then $\bigcup_{i\in\mathtt{support}(\sigma_{\mathtt{AG}})}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$. Then $\bigwedge_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\phi_{i}\wedge\neg\psi_{j}$ is in $\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})$. Note ${\sigma_{\mathrm{B}_{j}}}\subseteq{\sigma_{\mathtt{AG}}}$. Then $\bigwedge_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\phi_{i}\wedge\neg\psi_{j}$ is in $\mathtt{list}_{\mathrm{B}_{j}}({\sigma_{\mathrm{B}_{j}}})$. Clearly, $\bigwedge_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\phi_{i}\wedge\neg\psi_{j}\vDash_{\mathsf{MCL}+\mathsf{X}}\neg\psi_{j}$. Then $\neg\psi_{j}\in{\mathcal{E}_{\mathsf{X}}({\sigma_{\mathrm{B}_{j}}})}$.
    By Theorem 1, ${\sigma_{\mathrm{B}_{j}}}\ltimes_{(\mathtt{M},s)}\neg\psi_{j}$.
  • Assume $\mathsf{D}\in\mathsf{X}$. Consider whether $\mathrm{B}_{j}=\mathtt{AG}$.
    • Assume $\mathrm{B}_{j}=\mathtt{AG}$. Then $j\in\mathbb{J}_{0}$. Note $\mathtt{list}_{\mathtt{AG}}(\lambda_{\mathtt{AG}})\neq\emptyset$. Let $\chi\in\mathtt{list}_{\mathtt{AG}}(\lambda_{\mathtt{AG}})$. From the definition of $\mathtt{list}_{\mathtt{AG}}$, we can see $\neg\psi_{j}$ is a conjunct of $\chi$. Clearly, $\chi\vDash_{\mathsf{MCL}+\mathsf{X}}\neg\psi_{j}$. Note $\sigma_{\mathrm{B}_{j}}\subseteq\lambda_{\mathtt{AG}}$. Then $\chi\in\mathtt{list}_{\mathrm{B}_{j}}(\sigma_{\mathrm{B}_{j}})$. Then $\neg\psi_{j}\in{\mathcal{E}_{\mathsf{X}}({\sigma_{\mathrm{B}_{j}}})}$.
      By Theorem 1, ${\sigma_{\mathrm{B}_{j}}}\ltimes_{(\mathtt{M},s)}\neg\psi_{j}$.
    • Assume $\mathrm{B}_{j}\neq\mathtt{AG}$. Then there is $\overline{b}\in\mathtt{AG}-\mathrm{B}_{j}$. Let ${\sigma_{\mathtt{AG}}}\in\mathtt{JA}_{\mathtt{AG}}$ meeting the following conditions:
      1. (1)
        ${{\sigma_{\mathtt{AG}}}(b)}={{\sigma_{\mathrm{B}_{j}}}(b)}$ for all $b\in\mathrm{B}_{j}$;
      2. (2)
        ${{\sigma_{\mathtt{AG}}}(a)}=0$ for all $a\in\mathtt{AG}-\mathrm{B}_{j}$ such that $a\neq\overline{b}$.
      3. (3)
        ${{\sigma_{\mathtt{AG}}}(\overline{b})}=\begin{cases}j-\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})&\text{if }j\geq\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})\\ j-\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})+n&\text{otherwise}\end{cases}$
        where $n$ is the cardinality of $\mathbb{J}$.
      Note $j<n$. It is easy to see $\sigma_{\mathtt{AG}}(\overline{b})\in\mathbb{J}$. Clearly, ${\sigma_{\mathrm{B}_{j}}}\subseteq{\sigma_{\mathtt{AG}}}$.
      We claim $\mathtt{impeach}({\sigma_{\mathtt{AG}}})=j$.
      Note $\mathtt{impeach}({\sigma_{\mathtt{AG}}})=\sum\{\sigma_{\mathtt{AG}}(a)\in\mathbb{J}\mid a\in\mathtt{AG}\}\bmod n=\Big{(}\sum\{\sigma_{\mathtt{AG}}(a)\in\mathbb{J}\mid a\in\mathrm{B}_{j}\}+{{\sigma_{\mathtt{AG}}}(\overline{b})}\Big{)}\bmod n$. It suffices to show (($\bigg{(}\Big{(}\sum\{\sigma_{\mathtt{AG}}(a)\in\mathbb{J}\mid a\in\mathrm{B}_{j}\}\bmod n\Big{)}+\Big{(}{{\sigma_{\mathtt{AG}}}(\overline{b})}\bmod n\Big{)}\bigg{)}\bmod n=j$.
      Assume $j\geq\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})$. ($\Big{(}\sum\{\sigma_{\mathtt{AG}}(a)\in\mathbb{J}\mid a\in\mathrm{B}_{j}\}\bmod n\Big{)}+\Big{(}(j-\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}}))$ $\bmod n\Big{)}=\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})+(j-\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}}))=j$. $j\bmod n=j$. Thus, $\mathtt{impeach}({\sigma_{\mathtt{AG}}})$ $=j$.
      Assume $j<\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})$. ($\Big{(}\sum\{\sigma_{\mathtt{AG}}(a)\in\mathbb{J}\mid a\in\mathrm{B}_{j}\}\bmod n\Big{)}+\Big{(}(j-\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})+n)\bmod n\Big{)}=\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})+(j-\mathtt{impeach}({\sigma_{\mathrm{B}_{j}}})+n)=j+n$. $j+n\bmod n=j$. Then, $\mathtt{impeach}({\sigma_{\mathtt{AG}}})=j$.
      We claim $\mathtt{support}({\sigma_{\mathtt{AG}}})=\mathtt{support}({\sigma_{\mathrm{B}_{j}}})$. By Claim (2), $\mathtt{support}({\sigma_{\mathrm{B}_{j}}})\subseteq\mathtt{support}({\sigma_{\mathtt{AG}}})$. Let $i\in\mathtt{support}({\sigma_{\mathtt{AG}}})$. Then ${{\sigma_{\mathtt{AG}}}(a)}=i$ for all $a\in\mathrm{A}_{i}$.
        Assume $\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$. Note ${\sigma_{\mathrm{B}_{j}}}\subseteq{\sigma_{\mathtt{AG}}}$. Then ${{\sigma_{\mathrm{B}_{j}}}(a)}=i$ for all $a\in\mathrm{A}_{i}$. Then $i\in\mathtt{support}({\sigma_{\mathrm{B}_{j}}})$.
        Assume $\mathrm{A}_{i}\nsubseteq\mathrm{B}_{j}$. We want to show this case is impossible. Let $c\in\mathrm{A}_{i}-\mathrm{B}_{j}$. Then ${{\sigma_{\mathtt{AG}}}(c)}=i$. Note $c\in\mathtt{AG}-\mathrm{B}_{j}$. Then ${{\sigma_{\mathtt{AG}}}(c)}\in\mathbb{J}$, that is, $i\in\mathbb{J}$. We have a contradiction.
      Then $\mathtt{support}({\sigma_{\mathtt{AG}}})=\mathtt{support}({\sigma_{\mathrm{B}_{j}}})$.
      Note $\mathtt{support}({\sigma_{\mathrm{B}_{j}}})$ is $\mathsf{X}$-neat. Then $\mathtt{support}({\sigma_{\mathtt{AG}}})$ is $\mathsf{X}$-neat. Note $\bigcup_{i\in\mathtt{support}({\sigma_{\mathrm{B}_{j}}})}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$. Then $\bigcup_{i\in\mathtt{support}({\sigma_{\mathtt{AG}}})}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$. Therefore, $\bigwedge_{i\in\mathtt{support}({\sigma_{\mathrm{B}_{j}}})}\phi_{i}\wedge\neg\psi_{j}\wedge\bigwedge_{k\in\mathbb{J}_{0}}\neg\psi_{k}\in\mathtt{list}_{\mathtt{AG}}({\sigma_{\mathtt{AG}}})$. Note $\sigma_{\mathrm{B}_{j}}\subseteq\sigma_{\mathtt{AG}}$. Then $\bigwedge_{i\in\mathtt{support}({\sigma_{\mathrm{B}_{j}}})}\phi_{i}\wedge\neg\psi_{j}\wedge\bigwedge_{k\in\mathbb{J}_{0}}\neg\psi_{k}\in\mathtt{list}_{\mathrm{B}_{j}}({\sigma_{\mathrm{B}_{j}}})$. Clearly, $\bigwedge_{i\in\mathtt{support}({\sigma_{\mathrm{B}_{j}}})}\phi_{i}\wedge\neg\psi_{j}\wedge\bigwedge_{k\in\mathbb{J}_{0}}\neg\psi_{k}\models_{\mathsf{MCL}+\mathsf{X}}\neg\psi_{j}$. Then $\neg\psi_{j}\in{\mathcal{E}_{X}({\sigma_{\mathrm{B}_{j}}})}$.
      By Theorem 1, ${\sigma_{\mathrm{B}_{j}}}\ltimes_{(\mathtt{M},s)}\neg\psi_{j}$.
3.4 Upward derivability lemma
Lemma 3 (Upward derivability).
Let $\mathsf{X}\in\mathrm{ES}$ and $\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$ be a standard disjunction.
Then, $\vdash_{{\mathsf{MCL}+\mathsf{X}}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$, if the following condition, called the $\mathsf{X}$-derivability-reduction-condition, is met:
  1. (1)
    $\vdash_{{\mathsf{MCL}+\mathsf{X}}}\gamma$, or
  2. (2)
    the following two conditions hold:
    1. (a)
      if $\mathsf{D}\notin\mathsf{X}$, then there is $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$ such that $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat, $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\to\psi_{j}$;
    2. (b)
      if $\mathsf{D}\in\mathsf{X}$, then there is $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$ such that $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat, $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\to(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$.
Proof.
Assume $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma$. It is easy to see $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\to\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
Assume the following two conditions hold:
  1. (1)
    if $\mathsf{D}\notin\mathsf{X}$, then there is $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$ such that $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat, $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\to\psi_{j}$;
  2. (2)
    if $\mathsf{D}\in\mathsf{X}$, then there is $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$ such that $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat, $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\to(\psi_{j}\vee\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$.
We respectively consider whether $\mathsf{D}\in\mathsf{X}$.
  1. (1)
    Assume $\mathsf{D}\notin\mathsf{X}$. Then, there is $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$ such that $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat, $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\to\psi_{j}$. By Rule $\mathtt{Mon}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. We respectively consider whether $\mathsf{S}$ is in $\mathsf{X}$ and whether $\mathsf{I}$ is in $\mathsf{X}$.
    1. (a)
      Assume $\mathsf{S}\notin\mathsf{X}$ and $\mathsf{I}\notin\mathsf{X}$. Note in this case, $\mathbb{I}^{\prime}$ is $\epsilon$-neat: $\mathbb{I}^{\prime}$ is not empty, and for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $\mathrm{A}_{i}\neq\emptyset$ and $\mathrm{A}_{i^{\prime}}\neq\emptyset$, then $i=i^{\prime}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{SIA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
    2. (b)
      Assume $\mathsf{S}\notin\mathsf{X}$ and $\mathsf{I}\in\mathsf{X}$. Note in this case, Axiom $\mathtt{A}\text{-}\mathtt{IA}$ is an axiom of ${\mathsf{MCL}+\mathsf{X}}$, and $\mathbb{I}^{\prime}$ is $\mathsf{I}$-neat: $\mathbb{I}^{\prime}$ is not empty, and for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $i\neq i^{\prime}$, then $\mathrm{A}_{i}\cap\mathrm{A}_{i^{\prime}}=\emptyset$. By Axiom $\mathtt{A}\text{-}\mathtt{IA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
    3. (c)
      Assume $\mathsf{S}\in\mathsf{X}$ and $\mathsf{I}\notin\mathsf{X}$. Note in this case, $\mathtt{A}\text{-}\mathtt{Ser}$ is an axiom of ${\mathsf{MCL}+\mathsf{X}}$, and $\mathbb{I}^{\prime}$ is $\mathsf{S}$-neat: for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $\mathrm{A}_{i}\neq\emptyset$ and $\mathrm{A}_{i^{\prime}}\neq\emptyset$, then $i=i^{\prime}$.
      Assume $\mathbb{I}^{\prime}$ is empty. Note in this case, $\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle[\mathrm{B}_{j}]\rangle\psi_{j}$ is $\langle[\emptyset]\rangle\top\rightarrow\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\emptyset]\rangle\top\rightarrow\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, we have $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
      Assume $\mathbb{I}^{\prime}$ is not empty. By Axiom $\mathtt{A}\text{-}\mathtt{SIA}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$ $\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
    4. (d)
      Assume $\mathsf{S}\in\mathsf{X}$ and $\mathsf{I}\in\mathsf{X}$. Note in this case, Axioms $\mathtt{A}\text{-}\mathtt{Ser}$ and $\mathtt{A}\text{-}\mathtt{IA}$ are axioms of ${\mathsf{MCL}+\mathsf{X}}$, and $\mathbb{I}^{\prime}$ is $\mathsf{S}\mathsf{I}$-neat: for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $i\neq i^{\prime}$, then $\mathrm{A}_{i}\cap\mathrm{A}_{i^{\prime}}=\emptyset$.
      Assume $\mathbb{I}^{\prime}$ is empty. Note in this case, $\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle[\mathrm{B}_{j}]\rangle\psi_{j}$ is $\langle[\emptyset]\rangle\top\rightarrow\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\emptyset]\rangle\top\rightarrow\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, we have $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
      Assume $\mathbb{I}^{\prime}$ is not empty. By Axiom $\mathtt{A}\text{-}\mathtt{IA}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$ $\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
  2. (2)
    Assume $\mathsf{D}\in\mathsf{X}$. Then then there is $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$ such that $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat, $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\to(\psi_{j}\vee\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. By Rule $\mathtt{Mon}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. Note in this case, Axiom $\mathtt{A}\text{-}\mathtt{Det}$ is an axiom of ${\mathsf{MCL}+\mathsf{X}}$. We respectively consider whether $\mathsf{S}$ is in $\mathsf{X}$ and whether $\mathsf{I}$ is in $\mathsf{X}$.
    1. (a)
      Assume $\mathsf{S}\notin\mathsf{X}$ and $\mathsf{I}\notin\mathsf{X}$. Note in this case, $\mathbb{I}^{\prime}$ is $\mathsf{D}$-neat: $\mathbb{I}^{\prime}$ is not empty, and for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $\mathrm{A}_{i}\neq\emptyset$ and $\mathrm{A}_{i^{\prime}}\neq\emptyset$, then $i=i^{\prime}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{SIA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{Det}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})\rightarrow(\langle[\mathrm{B}_{j}]\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle[\mathtt{AG}]\rangle\psi_{k})$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}(\langle[\mathrm{B}_{j}]\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle[\mathtt{AG}]\rangle\psi_{k})\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
    2. (b)
      Assume $\mathsf{S}\notin\mathsf{X}$ and $\mathsf{I}\in\mathsf{X}$. Note in this case, Axiom $\mathtt{A}\text{-}\mathtt{IA}$ is an axiom of ${\mathsf{MCL}+\mathsf{X}}$, and $\mathbb{I}^{\prime}$ is $\mathsf{I}\mathsf{D}$-neat: $\mathbb{I}^{\prime}$ is not empty, and for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $i\neq i^{\prime}$, then $\mathrm{A}_{i}\cap\mathrm{A}_{i^{\prime}}=\emptyset$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{IA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{Det}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})\rightarrow(\langle[\mathrm{B}_{j}]\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle[\mathtt{AG}]\rangle\psi_{k})$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}(\langle[\mathrm{B}_{j}]\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle[\mathtt{AG}]\rangle\psi_{k})\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
    3. (c)
      Assume $\mathsf{S}\in\mathsf{X}$ and $\mathsf{I}\notin\mathsf{X}$. Note in this case, Axiom $\mathtt{A}\text{-}\mathtt{Ser}$ is an axiom of ${\mathsf{MCL}+\mathsf{X}}$, and $\mathbb{I}^{\prime}$ is $\mathsf{S}\mathsf{D}$-neat: for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $\mathrm{A}_{i}\neq\emptyset$ and $\mathrm{A}_{i^{\prime}}\neq\emptyset$, then $i=i^{\prime}$.
      Assume $\mathbb{I}^{\prime}$ is empty. Note in this case, $\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$ is $\langle[\emptyset]\rangle\top\rightarrow\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\emptyset]\rangle\top\rightarrow\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
      Assume $\mathbb{I}^{\prime}$ is not empty. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{SIA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{Det}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})\rightarrow(\langle[\mathrm{B}_{j}]\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle[\mathtt{AG}]\rangle\psi_{k})$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}(\langle[\mathrm{B}_{j}]\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle[\mathtt{AG}]\rangle\psi_{k})\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}$ $\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
    4. (d)
      Assume $\mathsf{S}\in\mathsf{X}$ and $\mathsf{I}\in\mathsf{X}$. Note in this case, Axioms $\mathtt{A}\text{-}\mathtt{Ser}$ and $\mathtt{A}\text{-}\mathtt{IA}$ are axioms of ${\mathsf{MCL}+\mathsf{X}}$, and $\mathbb{I}^{\prime}$ is $\mathsf{S}\mathsf{I}\mathsf{D}$-neat: for all $i,i^{\prime}\in\mathbb{I}^{\prime}$, if $i\neq i^{\prime}$, then $\mathrm{A}_{i}\cap\mathrm{A}_{i^{\prime}}=\emptyset$.
      Assume $\mathbb{I}^{\prime}$ is empty. Note in this case, $\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\rightarrow\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$ is $\langle[\emptyset]\rangle\top\rightarrow\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\emptyset]\rangle\top\rightarrow\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. By Axiom $\mathtt{A}\text{-}\mathtt{Ser}$, $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
      Assume $\mathbb{I}^{\prime}$ is not empty. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{IA}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\langle[\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}]\rangle\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}$. By repeated applications of Axiom $\mathtt{A}\text{-}\mathtt{Det}$, we can get $\vdash_{\mathsf{MCL}+\mathsf{X}}\langle[\mathrm{B}_{j}]\rangle(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})\rightarrow(\langle[\mathrm{B}_{j}]\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle[\mathtt{AG}]\rangle\psi_{k})$. Clearly, $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigwedge_{i\in\mathbb{I}^{\prime}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}(\langle[\mathrm{B}_{j}]\rangle\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\langle[\mathtt{AG}]\rangle\psi_{k})\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j}$. Then, $\vdash_{\mathsf{MCL}+\mathsf{X}}\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}$ $\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
3.5 Completeness by induction
Theorem 2 (Soundness and completeness of ${\mathsf{MCL}+\mathsf{X}}$).
For any $\mathsf{X}\in\mathrm{ES}$, the axiomatic system for ${\mathsf{MCL}+\mathsf{X}}$ given in Definition 7 is sound and complete with respect to the set of $\mathsf{X}$-valid formulas in $\Phi$.
Proof.
Let $\mathsf{X}\in\mathrm{ES}$. The soundness of the system for ${\mathsf{MCL}+\mathsf{X}}$ is easy to show, and we skip its proof. The completeness is shown as follows.
  1. (1)
    Let $\phi\in\Phi$. Assume $\models_{\mathsf{MCL}+\mathsf{X}}\phi$. We want to show $\vdash_{\mathsf{MCL}+\mathsf{X}}\phi$. We put an induction on the modal depth of $\phi$.
  2. (2)
    Assume the modal depth of $\phi$ is 00.
    The $\phi$ is a formula of the classical propositional logic. As ${\mathsf{MCL}+\mathsf{X}}$ extends the classical propositional logic, $\vdash_{\mathsf{MCL}+\mathsf{X}}\phi$.
  3. (3)
    Assume the modal depth of $\phi$ is $n$, which is greater than 00.
    By Lemma 1, there is $\phi^{\prime}$ such that (1) $\vdash_{\mathsf{MCL}+\mathsf{X}}\phi\leftrightarrow\phi^{\prime}$, (2) $\phi^{\prime}$ has the modal depth $n$, and (3) $\phi^{\prime}$ is in the form of $\chi_{0}\land\dots\land\chi_{k}$, where every $\chi_{i}$ is a standard disjunction. By soundness, $\vDash_{\mathsf{MCL}+\mathsf{X}}\phi\leftrightarrow\phi^{\prime}$. Then $\vDash_{\mathsf{MCL}+\mathsf{X}}\phi^{\prime}$. Let $i\leq k$. Then $\models_{\mathsf{MCL}+\mathsf{X}}\chi_{i}$. It suffices to show $\vdash_{\mathsf{MCL}+\mathsf{X}}\chi_{i}$.
  4. (4)
    Assume the modal degree of $\chi_{i}$ is less than $n$. By the inductive hypothesis, $\vdash_{\mathsf{MCL}+\mathsf{X}}\chi_{i}$.
  5. (5)
    Assume the modal degree of $\chi_{i}$ is $n$. Let $\chi_{i}=\gamma\lor(\bigwedge_{i\in\mathbb{I}}\langle[\mathrm{A}_{i}]\rangle\phi_{i}\rightarrow\bigvee_{j\in\mathbb{J}}\langle[\mathrm{B}_{j}]\rangle\psi_{j})$.
    By the downward validity lemma, the $\mathsf{X}$-validity-reduction-condition of $\chi_{i}$ is met:
    1. (a)
      $\models_{{\mathsf{MCL}+\mathsf{X}}}\gamma$, or
    2. (b)
      the following two conditions hold:
      1. i.
        if $\mathsf{D}\notin\mathsf{X}$, then there is $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$ such that $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat, $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, and $\models_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\to\psi_{j}$;
      2. ii.
        if $\mathsf{D}\in\mathsf{X}$, then there is $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$ such that $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat, $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, and $\models_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\to(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$.
    By the inductive hypothesis, the $\mathsf{X}$-derivability-reduction-condition of $\chi_{i}$ is met:
    1. (a)
      $\vdash_{{\mathsf{MCL}+\mathsf{X}}}\gamma$, or
    2. (b)
      the following two conditions hold:
      1. i.
        if $\mathsf{D}\notin\mathsf{X}$, then there is $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$ such that $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat, $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\to\psi_{j}$;
      2. ii.
        if $\mathsf{D}\in\mathsf{X}$, then there is $\mathbb{I}^{\prime}\subseteq\mathbb{I}$ and $j\in\mathbb{J}$ such that $\mathbb{I}^{\prime}$ is $\mathsf{X}$-neat, $\bigcup_{i\in\mathbb{I}^{\prime}}\mathrm{A}_{i}\subseteq\mathrm{B}_{j}$, and $\vdash_{\mathsf{MCL}+\mathsf{X}}\bigwedge_{i\in\mathbb{I}^{\prime}}\phi_{i}\to(\psi_{j}\lor\bigvee_{k\in\mathbb{J}_{0}}\psi_{k})$.
    By the upward derivability lemma, $\vdash_{\mathsf{MCL}+\mathsf{X}}\chi_{i}$.
4 Further work
There are two important topics for future work.
$\mathsf{CL}$ is decidable and the satisfiability problem for it is $\mathrm{PSPACE}$ [8]. By employing the downward validity lemma and the upward derivability lemma, it can be shown that the other seven coalition logics are also decidable. What are their computational complexities? We leave this as further work.
$\mathsf{ATL}$ is a temporal extension of $\mathsf{CL}$, whose featured formulas are as follows:
  • $\langle[\mathrm{A}]\rangle\mathsf{X}\phi$: some joint strategy of $\mathrm{A}$ ensures $\phi$ at the next moment.
  • $\langle[\mathrm{A}]\rangle\mathsf{G}\phi$: some joint strategy of $\mathrm{A}$ ensures $\phi$ at all moments in the future.
  • $\langle[\mathrm{A}]\rangle(\phi\mathsf{U}\psi)$: some joint strategy of $\mathrm{A}$ ensures $\phi$ until $\psi$.
$\mathsf{ATL}$ is a very useful logic. What are the logics which are temporal extensions of the other seven coalition logics? We will study this in the future.
Acknowledgments
Thanks go to the audience of seminars at University of Toulouse and University of Bergen, and the Third International Workshop on Logic of Multi-agent Systems at Zhejiang University, especially Thomas Ågotnes, Rustam Galimullin, and Emiliano Lorini.